OpenStudy (anonymous):

Determine the area of a square w/ the same perimeter as a regular hexagon w/ area 24square root3.

OpenStudy (anonymous):

P_hexagon = 6t P_square = 4s 6t = 4s s = 3t/2 A_hexagon = 3√(3)t²/2 = 24√3 t² = 16 t = 4 A_square = s² A_square = (3t/2)² A_square = (3*4/2)² A_square = 6² A_square = 36 http://en.wikipedia.org/wiki/Hexagon#Regular_hexagon

OpenStudy (anonymous):

Thank u so so so much took me a minute thanks tho!!!